Rx: Zephiran Chloride Solution (17% w/v) q.s. Purified Water, to make 480 mL Sig: one tbsp....

  1. Rx: Zephiran Chloride Solution (17% w/v) q.s.
    Purified Water, to make 480 mL
    Sig: one tbsp. diluted to 1 gallon with water to make a 1:10,000 dilution.

    How many milliliters of zephiran chloride solution should be used in preparing the prescription?
Answer #1
1 teaspoon = 5 ml 1 gallon= 4 L 1:10000 means 1 gm in 10000Ml, hence 0.1 gm in 1000 ml. This means 1 teaspoon (5 ml) of stock solution contain 0.1 gm zephiran chloride. Hence 480 ml = 9.6gm. 0.1/5* ( 480)= 9.6 gm. We need to know how many ml of 17% concentrate contains 9.6 gm. 17% w/v means 17 g in 100 ml Hence dose require will be (9.6/480)* 100= 56.4 ml
source: Books
answered by: Neethuben
Answer #2
1 teaspoon = 5 ml 1 gallon= 4 L 1:10000 means 1 gm in 10000Ml, hence 0.1 gm in 1000 ml. This means 1 teaspoon (5 ml) of stock solution contain 0.1 gm zephiran chloride. Hence 480 ml = 9.6gm. 0.1/5* ( 480)= 9.6 gm. We need to know how many ml of 17% concentrate contains 9.6 gm. 17% w/v means 17 g in 100 ml Hence dose require will be (9.6/480)* 100= 56.4 ml
source: Books
answered by: Neethuben
Answer #3
Pharmaceutical calculation
source: simoon
answered by: Elia